OpenStudy (anonymous):

What is the standard form of the number shown in the calculator display? 3.82e+7

OpenStudy (thomas5267):

\[ 3.82e{+7}=3.82\times10^7\\ \text{Another example:}\\ 5.34e{-3}=5.34\times10^{-3} \]

OpenStudy (anonymous):

i still dont get is

OpenStudy (anonymous):

this

OpenStudy (anonymous):

ok so i put it in my calculator and it says the answer is 38,200,000

OpenStudy (thomas5267):

\[ \color{green}{3.82}\color{cyan}{e}\color{red}{{+7}}=\color{green}{3.82}\times\color{cyan}{10}^\color{red}{7}\\ \text{Another example:}\\ \color{green}{5.34}\color{cyan}{e}\color{red}{{-3}}=\color{green}{5.34}\times\color{cyan}{10}^\color{red}{-3}\\ 3.82e{+7}=3.82\times10^7=3.82\times\underbrace{10,000,000}_{\text{7 zeros}} \]

OpenStudy (thomas5267):

=38,200,000

OpenStudy (anonymous):

so my calculator is right then ?

OpenStudy (anonymous):

and thx

OpenStudy (thomas5267):

Yes. A different expression of the same number.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

and here is ur medal

OpenStudy (thomas5267):

Thank you!

OpenStudy (anonymous):

welcome thx again

OpenStudy (thomas5267):

Unless your teacher specified to write it in this form, don't write numbers like this \( 5.34e{-3} \). Write this: \(5.34\times 10^{-3}\). Teachers don't like students to be lazy although they probably are!